OpenStudy (anonymous):

in preparation for the final exam, which counts as two exams, a student has scores of 88 and 97 and 91. what does the student need on the final exam to have a 90 average?

OpenStudy (anonymous):

(88+97+91+x)/4=90 360-88-97-91=x=84

OpenStudy (anonymous):

oh wait, thats also wrong! counts as two exams...

OpenStudy (anonymous):

(88+97+91+2x)/5=90 x=87

OpenStudy (anonymous):

just curious, what are these questions for?

OpenStudy (anonymous):

math final i am in a remedial math class

OpenStudy (anonymous):

I see... keep working at it!